OpenStudy (anonymous):

Line s passes through the points (2, 7) and (6, 1). Line t passes through the point (5, 2) and is perpendicular to line s. What is the equation for line t?

OpenStudy (anonymous):

To write the equation of a line, you need the slope, and either the y-intercept or a point. Since you have a point, we can write it in point slope form, which is (y-y1)=m(x-x1). But you'd still need to find the slope for line t. If the two lines are perpendicular, then the slope of line t is the negative reciprocal of the slope of line s. So the best way to do this (i think) would be Find the slope of line s using the two points given Find the reciprocal, and multiply times -1. This is the slope of line t, or m Then plug this into your point slope equation. You'll get y-2=m(x-5). You'll need to plug in m, of course. Once you finish that, you'll have an equation. If your equation needs to be in slope intercept form, which is y=mx+b, then take your equation and solve for y. Hope this helps!

OpenStudy (anonymous):

Thank you very much for you assistance. Your clarification has assured me on the solution.

OpenStudy (anonymous):

No problem, glad to help!